# Learning Fricke signs from Maass Form Coefficients

### v1

@@ Abstract.
KHL was partially supported by a grant from the Simons Foundation (#712100).
DLD was supported by the Simons Collaboration in Arithmetic Geometry, Number Theory, and Computation via the Simons Foundation grant 546235.
YDQ was supported by the NSF grant PHY-2019786 (the NSF AI Institute for Artificial Intelligence and Fundamental Interactions).
@@ 1. Introduction
The machine learning experiments presented here use the database of Maass forms from the LMFDB [15]. Among the 35,416 Maass forms in the LMFDB, some 15,423 of them lack rigorously computed Fricke signs, though it is possible to guess them heuristically. The computations of Maass forms underpinning the LMFDB rely on a combination of different techniques, including automorphic certification [3], explicit trace formulas [16], and an implementation of Hejhal’s algorithm [12] described in forthcoming work [13]. After that, the Fricke sign is subsequently determined by the signs of the coefficients $a_{p}$ for each prime $p$ dividing the level of the Maass form. In practice, however, it is often quite hard to compute the Maass form eigenvalue and coefficients with sufficient precision to directly recover the Fricke sign. This difficulty explains why a substantial portion of the dataset currently has unknown Fricke signs.
@@ 2.1. Definitions
For a broad overview, the reader is referred to [5]; a more extensive (but less approachable) reference is [4]. Let $\Delta$ denote the Laplace–Beltrami operator on the upper half-plane. A weight 00 Maass cuspform $f$ on $\Gamma_{0}(N)$ is a smooth square-integrable function $f:\mathcal{H}\rightarrow\mathbb{C}$ satisfying $f(\gamma z)=f(z)$, for all $\gamma\in\Gamma_{0}(N)$, and $(\Delta-\lambda)f(z)=0$, for some $\lambda\in\mathbb{C}$. We refer to $N$ as the level of $f$. Assuming the Selberg eigenvalue conjecture, we may write $\lambda=\frac{1}{4}+R^{2}$ for $R\in\mathbb{R}$. In what follows, we refer to $R$ as the spectral parameter.
@@
where $\epsilon$ is the root number. In particular, we have $\epsilon=(-1)^{\sigma(f)}w_{N}$ (see [4, §8]).
@@ 2.2. Averaging
Our analysis uses a dataset $\mathcal{L}$ containing the 35,416 rigorously computed Maass forms from the LMFDB [15]. The dataset contains the first $1,000$ Fourier coefficients $a_{n}$ for every Maass form, each of which has weight 00, trivial character, and integral level $N$ in the range from $1$ to $105$. In the analysis that follows, we are particularly interested in the parity $\sigma$ and the Fricke sign $w$ and (we will often drop the subscript from $w_{N}$ if there is no need to specify $N$). As discussed above, in the LMFDB, the value of the Fricke sign is not always rigorously computed, in which case it is given the value $w=0$. The breakdown of the Maass forms dataset $\mathcal{L}$ with different values for $\sigma$ and $w$ are given in Table 2.1.
@@ 2.3. Predicting the Fricke sign with LDA
The clear separation in Fricke sign provided by averaging $(-1)^{\sigma(f)}a_{p}$ indicates that it may be possible to predict the Fricke sign based on these features, perhaps even using a fairly simple technique. We choose to undertake a Linear Discriminant Analysis (LDA) to learn a linear decision boundary between classes of Fricke sign. LDA is a supervised dimensionality reduction and classification technique based on Bayes theorem [7, Section 4.3].
@@
Building on the paragraph above, we also make predictions using the trained LDA model for the Maass forms with unknown Fricke sign, that is, those in $\mathcal{L}_{0}$. Then, in order to check whether our predictions are reasonable, we examined if the average value of $(-1)^{\sigma(f)}a_{p}$ for the newly predicted Fricke sign follows similar murmuration patterns as the rigorously proved (known) values. Figures 2.3 and 2.4 demonstrate that our predicted values mimic the murmuration patterns seen in Figure 2.1. By comparing Figures 2.3 and 2.4 we make the interesting observation that, when $p$ is small, there is greater similarity between genuine and predicted murmurations for odd Maass forms than there is for even Maass forms.
@@ 2.4. Embedding the Fricke signs into the Fourier coefficients
$(a_{n})_{n=1}^{1000},~{}(a^{\prime}_{n})_{n=1}^{1000},~{}(a_{p_{i}})_{i=1}^{168},~{}(a^{\prime}_{p_{i}})_{i=1}^{168},$
where $p_{i}$ is the $i$th prime (there are $168$ primes less than $1000$). In these experiments, we see that replacing certain coefficients by zero does not make much difference to the accuracy. This is good news; we believe this means LDA should do well at predicting the Fricke sign in the case where it is unknown. Since the values of the coefficients are multiplicative, we would expect the $a_{p}$ to suffice. It is therefore surprising that there is a substantial improvement in our accuracy when using all $a_{n}$ rather than just $a_{p}$. We do not yet have a complete understanding of why this is true. It could be that the $a_{n}$ see the smaller primes more often due to their multiplicative nature, and these have a stronger impact on accuracy.
@@
We next analyze how these methods compare with each other when predicting the Fricke sign of Maass forms in $\mathcal{L}_{0}$. Predictions based on $(a_{n})_{n=1}^{1000}$ and $(a_{p_{i}})_{i=1}^{168}$ agree $83.29\%$ of the time. Predictions based on $(a_{n})_{n=1}^{1000}$ and $(a^{\prime}_{n})_{n=1}^{1000}$ agree $97.50\%$ of the time. Predictions based on $(a_{p_{i}})_{i=1}^{168}$ and $(a^{\prime}_{p_{i}})_{i=1}^{168}$ agree $99.81\%$ of the time. This is consistent with the accuracy results outlined above in that zeroing out the coefficients $a_{n}$ when $\gcd(n,N)>1$ does not have much impact on the predictions, but using $a_{n}$ versus $a_{p}$, does have a substantial impact on the predictions.
Finally, we analyze the different LDA methods used in this section and determine their accuracy relative to the number of coefficients used to train the data; see Figure 2.6. We again notice that LDA using $(a_{n})_{n=1}^{1000}$ has better accuracy than just using $(a_{p_{i}})_{i=1}^{168}$, and that there is little difference across the board for correct coefficients versus zeroed out coefficients. Interestingly, we note that there is striking improvement in accuracy for including more $a_{n}$ with small $n$, but then a leveling off as $n$ increases. This still occurs but is less pronounced when using $a_{p}$.
@@ 2.5. Predicting the Fricke sign with a neural network
$\tiny a_{2}$ $a_{3}$.. $a_{p_{d}}$ $R$$\tiny\mathsf{n}_{1}^{(1)}$ $\tiny\mathsf{n}_{2}^{(1)}$ $\tiny\mathsf{n}_{3}^{(1)}$ $\tiny\mathsf{n}_{4}^{(1)}$ $\tiny\mathsf{n}_{5}^{(1)}$ $\mathsf{n}_{1}^{(2)}$ $\mathrm{Prob}(w_{N}=1)$$\mathrm{ReLu}$$\sigma$

### v2

@@ 1. Introduction
The machine learning experiments presented here use the database of Maass forms from the LMFDB [18]. Among the 35,416 Maass forms in the LMFDB, some 15,423 of them lack rigorously computed Fricke signs, though it is possible to guess them heuristically. The computations of Maass forms underpinning the LMFDB rely on a combination of different techniques [15], including automorphic certification [4], explicit trace formulas [20], and an implementation of Hejhal’s algorithm [7] described in forthcoming work [16]. After that, the Fricke sign is subsequently determined by the signs of the coefficients $a_{p}$ for each prime $p$ dividing the level of the Maass form. In practice, however, it is often quite hard to compute the Maass form eigenvalue and coefficients with sufficient precision to directly recover the Fricke sign. This difficulty explains why a substantial portion of the dataset currently has unknown Fricke signs.
@@
Furthermore, the decomposition of the Fricke sign into a product of local factors implies that it may be viewed as a parity function in the sense of [21], where it is noted that general purpose machine learning methods have not previously succeeded in learning such functions from sign patterns. In Section 2.4, we will show that, although sign patterns are implicit in our data presentation, our classifiers are learning something more from the input features.
@@
Acknowledgments
Costa was supported by Simons Foundation grants 550033 and SFI-MPS-Infrastructure-00008651. Lee was supported by Simons Foundation grant 712100. Lowry-Duda was supported by Simons Foundation grant 546235. Qi was supported by the NSF grant PHY-2019786 (the NSF AI Institute for Artificial Intelligence and Fundamental Interactions).
We would also like to express our gratitude to the organizers of the Mathematics and Machine Learning Program at CMSA during the Fall 2024 semester, where this project was initiated.
@@ 2.1. Definitions
For a broad overview, the reader is referred to [6]; a more extensive (but less approachable) reference is [5]. Let $\Delta$ denote the Laplace–Beltrami operator on the upper half-plane. A weight 00 Maass cuspform $f$ on $\Gamma_{0}(N)$ is a smooth square-integrable function $f:\mathcal{H}\rightarrow\mathbb{C}$ satisfying $f(\gamma z)=f(z)$, for all $\gamma\in\Gamma_{0}(N)$, and $(\Delta-\lambda)f(z)=0$, for some $\lambda\in\mathbb{C}$. We refer to $N$ as the level of $f$. Assuming the Selberg eigenvalue conjecture, we may write $\lambda=\frac{1}{4}+R^{2}$ for $R\in\mathbb{R}_{\geq 0}$. In what follows, we refer to $R$ as the spectral parameter.
@@
where $\epsilon$ is the root number. In particular, we have $\epsilon=(-1)^{\sigma(f)}w_{N}$ (see [5, §8]).
@@ 2.2. Averaging
Our analysis uses a dataset $\mathcal{L}$ containing the 35,416 rigorously computed Maass forms from the LMFDB [18]. The dataset contains the first 1,000 Fourier coefficients $a_{n}$ for every Maass form, each of which has weight 00, trivial character, and integral level $N$ in the range from $1$ to $105$ [17]. In the analysis that follows, we are particularly interested in the parity $\sigma$ and the Fricke sign $w$ (we will often drop the subscript from $w_{N}$ if there is no need to specify $N$). As discussed above, in the LMFDB, the value of the Fricke sign is not always rigorously computed, in which case it is given the value $w=0$. The breakdown of the Maass forms dataset $\mathcal{L}$ with different values for $\sigma$ and $w$ are given in Table 2.1.
@@ 2.3. Predicting the Fricke sign with LDA
The clear separation in Fricke sign provided by averaging $(-1)^{\sigma(f)}a_{p}$ indicates that it may be possible to predict the Fricke sign based on these features, perhaps even using a fairly simple technique. We choose to undertake Linear Discriminant Analysis (LDA) to learn a linear decision boundary between classes of Fricke sign. LDA is a supervised dimensionality reduction and classification technique based on Bayes theorem [12, Section 4.3].
@@
LDA is likely to work well when the covariance is the same for all classes. To determine if LDA is a good technique for $\mathcal{D}$, we examined the covariance of $\mathcal{D}_{1}$ and $\mathcal{D}_{-1}$. Visual inspection of covariance matrices, their eigenvalues, and their determinants exhibited similar results between the two classes. To be more rigorous, we also applied Box’s M test [2] for each feature $a_{n}$. Without any further feature engineering, Box’s M test declared equal covariance for $641$ of the $1000$ features, including all prime indices. We note that the distribution of the features $a_{p}$ indexed by primes $p$ has spikes because the fixed value $a_{p}=-w_{p}/\sqrt{p}$ is repeated when $p$ divides the level (cf. Section 2.4). By way of example, in Figure 2.3, we plot the distribution of $a_{p}$ in the case that $p=7$, firstly over all forms in ±$\mathcal{L}_{\pm 1}$ and then only over those with level coprime to $7$.
Figure 2.3. Comparing the distribution of $a_{7}$ for Maass forms in $\mathcal{L}_{1}$ and $\mathcal{L}_{-1}$. In the left (resp. right) frame, we consider all levels (resp. only levels co-prime to $7$). The brown represents areas where the histograms overlap, and the green arc in the right frame is the semi-circle $y=\frac{1}{2\pi}\sqrt{4-x^{2}}$, given by the (vertical) Sato–Tate distribution [19].
With this in mind, for each $n$, we removed from $\mathcal{L}$ all Maass forms for which $n$ divides its level. Upon doing so, Box’s M test declared that the two classes had equal covariance for all but $33$ of the $1000$ features, all of which had composite index. Altogether, we are satisfied that the classes have equal covariance for the vast majority of the $n$. We further explore the impact of the values of $a_{n}$ for $n$ dividing the level in Section 2.4.
@@
Building on the paragraph above, we also make predictions using the trained LDA model for the Maass forms with unknown Fricke sign, that is, those in $\mathcal{L}_{0}$. Then, in order to check whether our predictions are reasonable, we examined if the average value of $(-1)^{\sigma(f)}a_{p}$ for the newly predicted Fricke sign follows similar murmuration patterns as the rigorously proved (known) values. Figures 2.4 and 2.5 demonstrate that our predicted values mimic the murmuration patterns seen in Figure 2.1. By comparing Figures 2.4 and 2.5 we make the interesting observation that, when $p$ is small, there is greater similarity between genuine and predicted murmurations for odd Maass forms than there is for even Maass forms.
@@ 2.4. Embedding the Fricke signs into the Fourier coefficients
$(a_{n})_{n=1}^{1000},~{}(a^{\prime}_{n})_{n=1}^{1000},~{}(a_{p})_{p<1000},~{}(a^{\prime}_{p})_{p<1000},$
where the indices range over the $168$ primes less than $1000$. In these experiments, we see that replacing certain coefficients by zero does not make much difference to the accuracy. This is good news; we believe this means LDA should do well at predicting the Fricke sign in the case where it is unknown. Since the values of the coefficients are multiplicative, we would expect the $a_{p}$ to suffice. It is therefore surprising that there is a substantial improvement in our accuracy when using all $a_{n}$ rather than just $a_{p}$. We will revisit this theme below in greater depth.
@@
We next analyze how these methods compare with each other when predicting the Fricke sign of Maass forms in $\mathcal{L}_{0}$. Predictions based on $(a_{n})_{n=1}^{1000}$ and $(a_{p})_{p<1000}$ agree $83.29\%$ of the time. Predictions based on $(a_{n})_{n=1}^{1000}$ and $(a^{\prime}_{n})_{n=1}^{1000}$ agree $97.50\%$ of the time. Predictions based on $(a_{p})_{p<1000}$ and $(a^{\prime}_{p})_{p<1000}$ agree $99.81\%$ of the time. This is consistent with the accuracy results outlined above in that zeroing out the coefficients $a_{n}$ when $\gcd(n,N)>1$ does not have much impact on the predictions, but using $a_{n}$ versus $a_{p}$ does have a substantial impact on the predictions.
Finally, we analyze the different LDA methods used in this section and determine their accuracy relative to the number of coefficients used to train the data; see Figure 2.7. We again notice that LDA using $(a_{n})_{n=1}^{1000}$ has better accuracy than just using $(a_{p})_{p<1000}$, and that there is little difference across the board for correct coefficients versus zeroed out coefficients. Interestingly, we note that there is striking improvement in accuracy for including more $a_{n}$ with small $n$, but then a leveling off as $n$ increases. This still occurs but is less pronounced when using $a_{p}$.
@@
As mentioned earlier, it is surprising that LDA performs better when using $(a_{n})_{n=1}^{1000}$ than when using just $(a_{p})_{p<1000}$. Indeed, the Fourier coefficients are multiplicative, and so there isn’t more information in the composite indices. We ran several tests to see if we could explain this observation. One might speculate that $a_{p}$ for small primes $p$ are likely to have a larger impact on classification. If that were the case, then perhaps the LDA performance improved with the inclusion of composite indices because the multiplicative property was accentuating the power of prediction coming from $a_{p}$ with small prime factors. To investigate this, we represented each Maass form by vectors of the form $(a_{n})_{n\in S}$, where
$S=\{n\leq 1000\;|\;n\text{ is divisible by }2,3\text{ or }5\}.$
Unexpectedly, LDA performed poorly with this presentation, with only $78.5\%$ accuracy, even though it included $733$ of the $1000$ features in $\mathcal{D}$. Also including all the primes in this range, that is, using vectors of the form $(a_{n})_{n\in S^{\prime}}$ where
$S^{\prime}=\{n\leq 1000\;|\;n\text{ is divisible by }2,3\text{ or }5\text{ or $n$ is prime}\},$
the accuracy increased to $89.8\%$ with $898$ of the $1000$ original features. While this is a mild improvement over the $86.2\%$ accuracy when using only prime indices, it is nowhere near the $96.1\%$ accuracy when using all indices $\leq 1000$, even though it contains $90\%$ of those values. It seems clear that extra information from $a_{p}$ with small primes $p$ is not improving the performance of LDA. Perhaps, LDA is learning something useful from the multiplicative property itself. The subset of indices we found with the best accuracy was
$S^{\prime\prime}=\{n\leq 1000\;|\;n\text{ is divisible by 1 or 2 prime factors}\}.$
For this presentation, LDA predicted Fricke signs with an accuracy of $95.3\%$ using $702$ of the 1000 features. A few other interesting results are shown in Table 2.3.
Feature indices LDA accuracy
$\{n\in\mathbb{Z}:1\leq n\leq 1000\}$ 96.1%
$\{n\text{ prime}:1\leq n\leq 1000\}$ 86.2%
$\{n\text{ $45$-smooth}:1\leq n\leq 1000\}$ 75.3%
$\{n\text{ even}:1\leq n\leq 1000\}$ 70.6%
$\{n\text{ odd}:1\leq n\leq 1000\}$ 93.4%
$\{n\in\mathbb{Z}:1\leq n\leq 500\}$ 93.3%
Table 2.3. Accuracy of LDA when predicting Fricke sign from features indexed by various subsets of integers. All of the subsets in the bottom portion of the table have about $500$ elements. In particular, smoothness bound of $45$ was chosen with that in mind.
2.5. Other LDA analysis
It is possible that the success of LDA could be explained by some subset of Maass forms that are especially easy to predict. For example, perhaps Maass forms of some levels are easier to predict than others. It is easy to check that the distribution of Fricke signs by level is generally even between $\mathcal{L}_{1}$ and $\mathcal{L}_{-1}$ (cf. Figure 2.8). The only exception is level $N=1$, where the Fricke sign is always $1$.
Figure 2.8. Comparing the distribution of Fricke signs by level. Unknown Fricke signs are denoted by 0.
We might expect forms on prime level to be easier to predict than forms on composite level, as the global Fricke sign is a product of local Fricke signs corresponding to each prime factor of the level. Further, the overall quality of the Maass forms approximations in the LMFDB is better for lower level than higher level. (This is caused by multiple compounding factors, including less separation between consecutive eigenvalues, forcing Heisenberg-Uncertainty tradeoffs and less well-behaved test functions in the trace formula used to compute the Maass forms). This leads the the general positive slope in Figure 2.9.
Figure 2.9. Comparing the percentages of known and unknown Fricke sign by level.
Building on this theme, we experimented with many different subsets of Maass forms of various levels and did not find any that were especially easy or hard to predict. The only exception was that if the subset was really small then it was, not surprisingly, hard to predict as there was not enough training data. These experiments used all the features and are summarized in Figure 2.10.
Figure 2.10. Exploring accuracy of LDA on subsets of Maass forms for given levels. The size of the dot corresponds to the size of the training set.
@@ 2.6. Predicting the Fricke sign with a neural network
$a_{2}$ $a_{3}$.. $a_{p_{d}}$ $R$$\mathsf{n}_{1}^{(1)}$ $\mathsf{n}_{2}^{(1)}$ $\mathsf{n}_{3}^{(1)}$ $\mathsf{n}_{4}^{(1)}$ $\mathsf{n}_{5}^{(1)}$ $\mathsf{n}_{1}^{(2)}$ $\mathrm{Prob}(w_{N}=1)$$\mathrm{ReLu}$$\sigma$